OpenStudy (anonymous):

how can you find the polar coordinates of (-2, 5) & satisfy 0≤theta≤2pi and -pi ≤theta≤pi ?

OpenStudy (anonymous):

r=sqrt(2^2+5^2)=sqrt(29) theta=tan^-1(5/-2) x=r cos theta, y= r sin theta

OpenStudy (anonymous):

theta=292 for first case and theta=-68 ..........2 nd case

OpenStudy (anonymous):

oh! got it!

OpenStudy (anonymous):

in polar co -ordinates we wrie the points as (r, theta) for first case (sqrt(29), 292) for 2nd case (sqrt(29),-68)